# Folded multistability and hidden critical point in microwave-driven Rydberg atoms

### v2

@@
Methods
Details of the experimental setup
@@ Methods
Lindblad master equation
@@
We consider the loss of the microwave field and make use of the substitution relationship between microwave power $P$ and Rabi frequency $\Omega_{m}$, given by the formula $\Omega_{m}=10^{\frac{P}{20}}$. This allows us to obtain the dashed lines in Fig. 2(b) for the phase diagram of the system. By applying both forward and backward scans on $\Delta_{1}$, we can generate a hysteresis loop represented by the blue and red lines. In this process, we select the closest steady state at $\rho_{r_{1}r_{1}}[t+\Delta t]$ as the next step in the evolution of $\rho_{r_{1}r_{1}}[t]$. Consequently, this approach enables us to construct the hysteresis loop effectively.
Data Availability
All experimental data used in this study are available from the corresponding author upon request.
Acknowledgements
D.-S.D. thanked the discussions with Dr. Zong-Kai Liu. We acknowledge funding from the National Key R and D Program of China (Grant No. 2022YFA1404002), the National Natural Science Foundation of China (Grant Nos. U20A20218, 61525504, and 61435011), and the Major Science and Technology Projects in Anhui Province (Grant No. 202203a13010001).
Author contributions statement
D.-S.D. conceived the idea for the study. Y.M. conducted the physical experiments with B.L. and L.-H.Z. D.-S.D. and Y.J.W developed the theoretical model. The manuscript was written by D.-S.D. The research was supervised by D.-S.D. All authors contributed to discussions regarding the results and the analysis contained in the manuscript.
Competing interests
The authors declare no competing interests.

### v3

@@ VI Discussions
Acknowledgements
We acknowledge funding from the National Key R and D Program of China (Grant No. 2022YFA1404002), the National Natural Science Foundation of China (Grant Nos. U20A20218, 61525504, and 61435011), and the Major Science and Technology Projects in Anhui Province (Grant No. 202203a13010001). Y.M. conducted the physical experiments with B.L. and L.-H.Z. D.-S.D. and Y.J.W developed the theoretical model. The manuscript was written by D.-S.D. All authors contributed to discussions regarding the results and the analysis contained in the manuscript. D.-S.D. conceived the idea and supported this project.
APPENDIX A: Details of the experimental setup
@@ Acknowledgements
APPENDIX B: Lindblad master equation
@@
We consider the loss of the microwave field and use the substitution relationship between microwave power $P$ and Rabi frequency $\Omega_{m}$, given by the formula $\Omega_{m}=\frac{\mu}{\hbar\sqrt{\varepsilon_{0}cA}}10^{\frac{P}{20}-1.5}$ in the theoretical part, where $A$ represents the square of the region where Rydberg atoms interact with the microwave. This allows us to obtain the dashed lines in Fig. 2(b) for the phase diagram of the system. The relationship between output microwave power on the signal generator settings and the actual microwave power at the atoms has been measured, presenting a loss of 12.91 dBm from the signal generator to the atoms. By applying both forward and backward scans on $\Delta_{1}$, we can generate a hysteresis loop represented by the blue and red lines. In this process, we select the closest steady state at $\rho_{r_{1}r_{1}}[t+\Delta t]$ as the next step in the evolution of $\rho_{r_{1}r_{1}}[t]$. Consequently, this approach enables us to construct the hysteresis loop effectively.